OpenStudy (anonymous):

1/{(-2)^-2} + {(2)^-2} Please expand the above by first writing the exponents as positive. Then solve and show the answer as one simple fraction (the one without exponent)

OpenStudy (anonymous):

2^-1 = 1/2 so 2^-2 = 1/4 u can finish it

OpenStudy (anonymous):

Thank you so much!